Mr. Jones owns a small glass souvenir business. He hires one assistant for $ 12,000 a year, payable at the end of the year, and $ 20,000 a year to buy raw materials and supplies, payable at the beginning of the year. At the beginning of the year, for the purchase of new equipment, which has a lifespan of 8 years, Mr. Jones took a loan from the bank in the amount of 40 thousand dollars at 10% per annum with payment at the end of the year. The interest on deposits is equal to the interest on loans. Mr. Jones uses his own space as a workshop. He could rent it out for 10 thousand. dollars per year payable at the end of the year. Mr. Jones's competitor offers him a job as a glassblower at $ 15,000 a year, payable at the end of the year. The total annual income from the sale of souvenirs is 65 thousand dollars.

Find Mr. Jones' Accounting and Economic Profits.

 



1
Expert's answer
2021-02-24T15:03:45-0500
"TC=12000+20000+5500=37500"

"TR=65000"

Accounting


"TR-TC=27500"

Economic Profits


"27500-10000-15000=2500"